OpenStudy (anonymous):

A pound of grass seed covers 120 square feet of lawn. How many pounds are needed to seed a lawn measuring 50 by 40 feet (2000 sq ft)?

OpenStudy (anonymous):

IS IT 2000/120=16

OpenStudy (anonymous):

So we need 1 pound of grass seed for every 120 square feet of lawn. We have a total of 2000 sq feet of lawn to seed. Since we use 1 pound/120 sq ft, dividing the total amount of lawn by the 120 sq ft will give us the number of pounds of grass seed we need. \[\bf \frac{ 2000 \ \cancel{sq.ft} }{ 120 \ \cancel{sq.ft} }=[ \ \ \ ] \ lbs.\]What goes inside the brackets? @wheelchairmomof2

OpenStudy (anonymous):

17

OpenStudy (anonymous):

Correct. So we need 17 lbs of grass seed to cover a 2000 sq.ft lawn. @wheelchairmomof2

OpenStudy (texaschic101):

proportion time... 1/120 = x/2000 (1 lb to 120 sq ft = x lbs to 2000 sq ft) cross multiply (120)(x) = (1)(2000) 120x = 2000 x = 2000/120 x = 16.6 sq ft

OpenStudy (anonymous):

@wheelchairmomof2 Why did you round it off to 17, why not keep at as 16.6? @texaschic101 Question asks for pounds of grass seed, not sq.ft of grass seed.

OpenStudy (texaschic101):

are you sure you would put 17 because 17 x 120 = 2040

OpenStudy (texaschic101):

my bad....lbs of grass seed

OpenStudy (anonymous):

16.6

OpenStudy (texaschic101):

thats what I would put

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ty

OpenStudy (texaschic101):

no problem :)